A. Aggregate Planning

Problem 1. AUTO Inc. are in the process of developing an aggregate plan for their latest sports car, model "Sporty-Sprinty". Anticipated sales forecast in A.U. for the next four months is tabulated below:

Month                         Jan                  Feb                Mar                 April

Sales forecast (A.U.)    100                   200                 300                   600

           

  • Initial inventory (at the beginning of first month) = 100 A.U.
  • Initial workforce level = 50
  • Minimum ending inventory required at the end of the plan = 30 A.U.
  • Minimum ending inventory needed at the end of any month other than the last = 0
  • AUTO Inc,. use 1 Worker-Month production (Production rate) = 10 a.u.
  • Salary of a regular time worker = $7,000 per month.
  • Hiring Cost per worker = $1,200
  • Layoff Cost per worker = $2,500
  • Inventory carrying cost per A.U. per month = $50
  • Overtime = 20% of regular time for two months only paid @ 150% of the regular time production rate.

AUTO Inc. do not allow shortages in their aggregate plan. Develop

  1. Level workforce with overtime plan,

(ii) Chase plan

Based on cost criterion which strategy would you recommend?
Answer #1
Month J F M A
Sales forecast 100 200 300 600
Production 300 300 300 300
Initial inventory 100 300 400 400
Ending inventory 300 400 400 100
Overtime 0 0 0 0
Hiring 0 0 0 0
Firing 20 0 0 0

Regular workers = 30

firing cost = 20*2500=50000

Workers' wages = 7000x30x4 = 840000

Inventory cost = avg inventory * cost of holding one unit per month

= 1200*50 =60000

Total cost = 50000+840000+60000 =950000

2. Chase plan

Month J F M A
Sales forecast 100 200 300 600
Production 0 200 300 630
Initial inventory 100 0 0 0
Ending inventory 0 0 0 30
Overtime 0 0 0 0
Hiring 0 20 10 33
Firing 50 0 0 0

Hiring cost =1200x63= 75600

Firing cost = 2500x50 =125000

Inventory cost = 50x65 =3250 ( based on avg inventory)

workers' wages = (20+30+63)x7000 =791000

Total cost = 75600+125000+3250+791000 =994850

It can be seen that level strategy is better than chase strategy as it has lower total cost.